OpenStudy (anonymous):

how do you get the inverse of this function? f(x)= 1/x+2

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

Set \[x = \frac{1}{y+2} \] and solve for y

OpenStudy (anonymous):

get \[xy + 2x = 1\]

OpenStudy (anonymous):

so \[y = \frac{1-2x}{x} = \frac{1}{x} - 2 \]

OpenStudy (anonymous):

that is the inverse function

OpenStudy (anonymous):

but how do you simplify the 1/y+2? sorry, i'm so bad at math.

OpenStudy (anonymous):

Well, the original function was \[\frac{1}{x+2}\] If you take the reciprocal of this you get \[x+2\] then if you subtract 2 you get \[x\], which is the identity function

OpenStudy (anonymous):

so the inverse function is "take reciprocal and subtract 2"

OpenStudy (anonymous):

that is , \[ \frac{1}{x} - 2\]

OpenStudy (anonymous):

ah that makes complete sense. thanks so much!!!

OpenStudy (anonymous):

No problem :) good luck with the rest of your work

OpenStudy (anonymous):

oh thansk! actually, one last question; when solving for inverses what are the usual steps you take? do you just switch x and y and simplify from there or...?